4. A fluid system undergoes a non-flow frictionless process from 4.5 ft 3 to a final volume 1.5 ft 3 in

accordance with the defining relation P= 60/V + 30 psi, where volume in ft 3 . During the process the

system reject 30 BTU of heat. Determine the change in enthalpy and internal energy.


1
Expert's answer
2021-09-16T13:08:49-0400

Given:

"V_1 = 4.5\; ft^3 \\

V_2 = 1.5 \; ft^3 \\

Q_{1-2} = - 30 \; BTU \\

P = \frac{60}{V} + 30 \; psi"

we know that work done

"W_{1-2} = \int^{V_2}_{V_1} PdV = \int^{V_2}_{V_1} [\frac{60}{V} +30]dV \\

W_{1-2} = \int^{V_2}_{V_1} \frac{60}{V} dV + \int^{V_2}_{V_1} 30dV \\

W_{1-2} = 60[ln(V)]^{V_2}_{V_1} + 30[V]^{V_2}_{ V_1} \\

W_{1-2} = 60 ln(\frac{V_2}{V_1}) + 30(V_2-V_1) \\

W_{1-2} = 60ln(\frac{1.5}{4.5}) + 30(1.5 - 4.5) \\

W_{1-2} = -155.9167 \; psi \cdot ft^3"

we know that

"1 \; psi = 144 \; lbf/ft^2 \\

1 \; BTU = 778 \; ft \cdot lbf \\

W_{1-2} = \frac{-155.9167 \times 144}{ 778} \; BTU \\

W_{1-2} = -28.86 \; BTU"

From first law of thermodynamics

"Q = ΔU + W \\

ΔU = Q - W \\

ΔU = -30 - (-28.86)"

ΔU = -1.14 BTU Change in inlernal energy

Now change in enthalpy is given as

"ΔH = ΔU + (P_2V_2 -P_1V_2) \\

P_1=\frac{60}{V_1} + 30 = \frac{60}{4.5} + 30 + 43.33 \; psi \\

P_1V_1=43.33 \times 4.5=195\; psi \cdot ft^3= \frac{195 \times 144}{778} = 36.09 \; BTU \\

P_2 = \frac{60}{V_2} + 30 = \frac{60}{1.5} + 30 = 70 \; psi \\

P_2V_2 = 70 \times 1.5 = 105 \; psi \cdot ft^3 = \frac{105 \times 144}{778} = 19.43 \; BTU

ΔH = ΔU + (P_2V_2 – P_1V_1) \\

= - 1.14 + (19.43 -36.09) \\

ΔH = - 17.8 \;BTU"

change in enthalpy